OpenStudy (anonymous):

Simplify the square root of -121

OpenStudy (anonymous):

11i

OpenStudy (mathstudent55):

\(\sqrt{-121} \) Remember that when you deal with the square root of a negative number, the answer is imaginary, and you need to use \(\sqrt{-1} = i\). \(\sqrt{-121} = \sqrt{-1 \times 121} = \sqrt{-1} \times \sqrt{121} = i \times 11 = 11i\)

OpenStudy (mathmale):

Hint: You must recognize that this is the square root of a negative number, and that Sqrt(-1) is defined to be a certain quantity. What is that quantity? What is the Sqrt of +121?

OpenStudy (anonymous):

thank you guys!